OpenStudy (anonymous):

Factor first, then simplify the factors: 2(x+4)(2x-1)^3 + 6(x+4)^2 (2x-1)^2

OpenStudy (anonymous):

\[2(x+4)(2x-1)^3 + 6(x+4)^2(2x-1)^2 \]

OpenStudy (anonymous):

How do I begin?

OpenStudy (anonymous):

Notice that there are 2 (x+4) term and 2 (2x-1) terms

OpenStudy (anonymous):

Or, if you REALLY need the solution, here it is. (WARNING: MAY NOT BE ACCURATE.) 2(x+4)(2x-1)^3 + 6(x+4)^2 (2x-1)^2 =2(x+4)(2x-1)^2((2x-1)+3(x+4)) =2(x+4)(2x-1)(2x-1+3x+12) =2(x+4)(2x-1)(5x-11) There you have it.

OpenStudy (anonymous):

find HCF of 2 &6. take the two factors (2x-1) and (x+4) along with their smallest power respectively.

OpenStudy (anonymous):

Alright. Thank you!